### v1

@@ Abstract
AI methods, such as generative models and reinforcement learning, have recently been applied to combinatorial optimization (CO) problems, especially NP-hard ones. This paper compares such GPU-based methods with classical CPU-based methods on Maximum Independent Set (MIS). Experiments on standard graph families show that AI-based algorithms fail to outperform and, in many cases, to match the solution quality of the state-of-art classical solver KaMIS running on a single CPU. Some GPU-based methods even perform similarly to the simplest heuristic, degree-based greedy. Even with post-processing techniques like local search, AI-based methods still perform worse than CPU-based solvers.
We develop a new mode of analysis to reveal that non-backtracking AI methods, e.g. LTFT (which is based on GFlowNets), end up reasoning similarly to the simplest degree-based greedy approach, and thus worse than KaMIS. We also find that CPU-based algorithms, notably KaMIS, have strong performance on sparse random graphs, which appears to refute a well-known conjectured upper bound for efficient algorithms from Coja-Oghlan & Efthymiou (2015).
@@ 1 Introduction
Combinatorial optimization (CO) lies at the core of numerous scientific and engineering studies, encompassing applications in network design, resource allocation, healthcare, and supply chain (Du & Pardalos, 1998; Hoffman, 2000; Zhong & Tang, 2021). Combinatorial optimization usually involves selecting an optimal solution from a discrete but often exponentially large set of candidates. Many are NP-hard (meaning that if $P\neq NP$ then there is no polynomial-time algorithm for solving them in the general cases (Papadimitriou & Steiglitz, 1998)). This makes it challenging to design algorithms with provable guarantees, but in practice solvers can find reasonable quality solutions (e.g., Gurobi (Gurobi Optimization, LLC, 2024)).
Recent advances in artificial intelligence (AI) and GPU computing have motivated use of AI-inspired approaches, e.g., Graph Neural Networs (GNNs) and reinforcement learning, to learn problem-specific strategies for NP-hard optimization problems such as MIS (Li et al., 2018; Ahn et al., 2020) and TSP (Kool et al., 2018; Zhang et al., 2021). AI models can also be trained to predict search directions or refine heuristic rules (Li et al., 2018; d O Costa et al., 2020). These algorithms utilize advanced GPUs and often require shorter inference compared to CPU-based algorithms. Additionally, AI-inspired methods avoid the need to design heuristics for specific problems, allowing generalization to new instances and problems (Bengio et al., 2021b; Cappart et al., 2023).
Despite the claimed benefits of AI-inspired methods, a few years ago Angelini & Ricci-Tersenghi (2023) showed that a Graph Neural Network based MIS algorithm (Schuetz et al., 2022) failed to surpass greedy algorithms. Böther et al. (2022) showed that AI-inspired approaches fail to provide superior search directions compared to traditional heuristics in tree search algorithms for MIS. Gamarnik (2023) suggests that GNN has theoretical limits which may become obstacles for GNN-based MIS algorithms.
However, in recent years many new AI-inspired CO algorithms, utilizing a variety of techniques, such as diffusion models (Sun & Yang, 2023; Sanokowski et al., 2024), GPU-accelerated sampling (Sun et al., 2023), and direct optimization (Alkhouri et al., 2024) have been developed and claimed to significant improve over the previous ones. Furthermore, GFlowNets Bengio et al. (2021a), which have been proposed as general-purpose tools for tasks like scientific discovery and reinforcement learning Bengio et al. (2023), has also been used to solve CO problems (Zhang et al., 2023). Since combinatorial optimization is an arena where humans have hand-designed algorithms for many decades, the following question is of great scientific interest:
Do AI-inspired algorithms perform better than classical heuristics for
combinatorial optimization?
1.1 Our contributions
We explore the question in the context of Maximum Independent Set (MIS) problem: given a graph, aiming to find the largest subset of nodes with no edges present between any node pair. The simplicity of the problem attracted design of many heuristics to tackle the problem (Andrade et al., 2012; Lamm et al., 2015). In recent works, MIS is also a main target in efforts that design AI-inspired approaches such as non-convex optimization (Schuetz et al., 2022; Alkhouri et al., 2024), reinforcement learning (Ahn et al., 2020; Zhang et al., 2023), and diffusion models (Sun & Yang, 2023; Sanokowski et al., 2024).
For classical heuristics, we test degree-based greedy (Deg-Greedy, pick a node with smallest degree at each step), and the state-of-the-art MIS solver KaMIS (Lamm et al., 2017; Dahlum et al., 2016; Hespe et al., 2019). For AI-inspired algorithms, we test the newest algorithms from each “category” according to the techniques they use, including sampling algorithm iSCO (Sun et al., 2023), non-convex optimization algorithm PCQO (Alkhouri et al., 2024), reinforcement-learning related algorithms LwD (Ahn et al., 2020) and LTFT (Zhang et al., 2023) (based on GFlowNets (Bengio et al., 2021a)) and diffusion models DIFUSCO (Sun & Yang, 2023) and DiffUCO (Sanokowski et al., 2024).
Testing on different graph types with different sizes and densities leads to the following empirical finding (Section 3).
Finding 1: Current AI-inspired algorithms for MIS still don’t outperform the best heuristic KaMIS, which runs on a single thread in a CPU, while AI-inspired methods often require significant computational resources.
Finding 2: As the graph becomes larger or denser, KaMIS exhibits a notable superiority to AI-inspired algorithms.
Finding 3: The simplest degree-based greedy algorithm (Deg-Greedy) serves as a very strong baseline. Some AI-inspired algorithms perform similarly to or worse than Deg-Greedy, especially for larger and denser graphs.
Section 4 presents ablation studies to understand why some AI-inspired methods fail to improve over the simplest Deg-Greedy method. We introduce a new mode of analysis, serialization, that transforms the solution of any algorithm into a sequential list of choices leading to the final independent set. We compare sequential order with the one produced by Deg-Greedy (Sections 4.1 and 4.2). We find that the reinforcement learning related algorithm LTFT, based on GFlowNets, behaves similarly to Deg-Greedy. We also found several qualitative characteristics that appear to distinguish algorithms that perform significantly better than Deg-Greedy from those that perform similarly or even worse than Deg-Greedy. We also explore whether AI-inspired method can improve their solution quality via a post-processing step using local search (Section 4.3), but find that they still fail to outperform KaMIS after some improvements. In Section 4.4, we discuss an additional result that may be of interest for MIS experts: on random graphs, KaMIS has a level of performance on MIS that appears to contradict a well-known conjecture about polynomial-time algorithms Coja-Oghlan & Efthymiou (2015).
2 Benchmarking MIS Algorithms
We focus the experiment setup for benchmarking different algorithms for Maximum Independent Set problems (MIS).
@@ 2.2 MIS algorithms
Deg-Greedy
(Degree-based greedy) Simplest heuristic: always picks a node with the smallest degree in the current graph, add to the current independent set, and delete that node and all of its neighbors from the graph. Most papers on AI-inspired methods do not compare with this baseline.
OnlineMIS and ReduMIS
are two variants of the MIS solver KaMIS, mainly consists of three alternating steps: greedy, local search, and graph reductions. OnlineMIS (Dahlum et al., 2016) only applies a simple reduction after local search, while ReduMIS Lamm et al. (2017) applies many graph reduction techniques.
iSCO
(Sun et al., 2023) is a GPU-accelerated sampling-based method, incorporating gradient-based discrete MCMC and simulated annealing. The MCMC is designed based on the Metropolis-Hasting algorithm, which if given enough time (exponential), can get the optimal solution.
PCQO
(Alkhouri et al., 2024) directly optimizes the quadratic loss function of the MIS using gradient descent. It is sensitive to optimization hyperparameters, so hyperparameter search is required for achieving good results.
LwD
(Ahn et al., 2020) is a reinforcement learning based algorithm which models the problem as a Markov Decision Process (MDP) and requires a dataset (without solutions) to train the policy. In each step, several nodes are added to the independent set and are never deleted. We call it a non-backtracking algorithm. Ahn et al. (2020) reported that it outperforms KaMIS on very large but very sparse random graphs, which we do not include in our benchmark.
LTFT
(Zhang et al., 2023) is also a non-backtracking MDP-based algorithm similar to LwD, but it only selects one node at a time, which is decided by GFlowNets (Bengio et al., 2021a). Thus, it has a very similar procedure to Deg-Greedy. The algorithm requires a dataset (without solutions) to train the neural network.
DIFUSCO and DiffUCO
DIFUSCO (Sun & Yang, 2023) is an end-to-end one-shot MIS solver using diffusion models and requires a supervised training dataset with solutions. DiffUCO (Sanokowski et al., 2024) also uses diffusion model but with unsupervised learning and annealing techniques. It requires a training dataset without solutions.
@@
CPU-based GPU-acc Learning-based $n$ $d$ Deg-Greedy OnlineMIS ReduMIS iSCO PCQO LwD LTFT DIFUSCO DiffUCO 100 10 29.25 30.50 30.50 30.62 25.00 30.38 28.62 30.25 30.02 30 13.63 14.00 14.75 14.50 11.50 14.38 13.12 13.88 13.92 300 10 77.50 93.88 94.38 94.75 90.62 94.25 88.62 93.50 91.84 30 44.50 47.88 47.88 47.62 41.75 46.88 43.25 43.88 45.04 100 16.13 18.00 18.38 18.00 13.38 17.00 16.25 16.62 16.93 1000 10 303.25 314.75 316.13 315.62 305.12 311.25 297.00 303.88 311.67 30 151.00 158.88 163.75 163.50 143.00 158.38 150.00 143.75 152.55† 100 60.63 64.75 66.63 66.50 55.62 63.88 60.88 55.38 63.55† 300 22.25 25.00 25.75 24.62 20.00 19.12* 22.62 20.88 18.36† 3000 10 907.13 947.25 954.25 950.88 919.00 934.12 888.25 902.00 935.28† 30 451.88 480.88 493.13 491.62 442.38 473.25 449.00 413.38 459.57† 100 183.63 194.38 201.50 200.38 170.38 190.75* 184.00 171.38 194.07† 300 73.50 77.63 80.75 78.88 63.50 73.88 1000 23.38 26.00 26.25 19.88 23.62 10000 10 2999.88 3161.88 3173.62 3149.92 2685.62 30 1498.00 1607.50 1639.88 1625.47 1334.75 100 613.75 650.00 670.88 573.63 300 249.00 258.38 272.25 239.13 1000 87.13 91.50 94.25 80.88 3000 29.63 33.00 33.25 19.00
2.3 Graph types
Erdős-Renyí (ER) graph
(Erdös & Rényi, 1959) are random graphs where edges are connected uniformly at random (with a given probability or a fixed number of edges). We vary $2$ parameters for ER graphs, number of nodes $n$ and average degree $d$, by fixing the number of edges at $\frac{nd}{2}$. Previous benchmark (Böther et al., 2022) and algorithms (Ahn et al., 2020; Sun & Yang, 2023; Zhang et al., 2023; Alkhouri et al., 2024) used it as test graphs for MIS, though without varying parameters as we did.
Barabási–Albert (BA) graph
(Albert & Barabási, 2002) are random graphs generated by a probabilistic growth process, mimicking real-world networks such as Internet, citation networks, and social networks (Albert & Barabási, 2002; Radicchi et al., 2011). For BA graphs, we vary $2$ parameters: number of nodes $n$ and parameter $m$ (not number of edges). The average degree of BA graphs can be approximated as $2m$.
Real-world graphs
We pick REDDIT-MULTI-5K and COLLAB (Yanardag & Vishwanathan, 2015) from TUDataset website (Morris et al., 2020), since they have enough graphs for training and graph sizes not too small. REDDIT-MULTI-5K has $508.52$ average nodes and $594.87$ average edges. They are mostly very sparse graphs. COLLAB has $74.49$ average nodes and 2457.78 average edges. They are mostly small but dense graphs.
2.4 More experiment details
We set $24$-hr time limit for KaMIS (OnlineMIS and ReduMIS) since it runs on a single CPU thread with 32GB memory, and our benchmark focuses on performance on solution quality. Note that AI-based methods run well only on relatively small graphs, and ReduMIS runs in less than one hour on small graphs $(n\leq 3000)$. Given 24 hours ReduMIS can handle much larger graphs (up to $n\approx 1e6$).
For iSCO and learning-based algorithms, we report results within our computational limit (generally a single 80GB A100 for $96$hrs, details in Appendix B.) We test PCQO for $n\leq 10000$ because the performance degrades quickly for large graphs using default hyperparameter search domain.
3 Experiment Results and Main Findings
In this section, we present our main experiment results. The performance of different algorithms on Erdős-Renyí (ER) graphs, Barabási–Albert (BA) graphs, and real-world graphs are shown in Tables 1, 2 and 3 respectively.
@@ 3 Experiment Results and Main Findings
Our first main finding is that, current AI-inspired algorithms do not outperform the best classical heuristics ReduMIS in terms of performance. As shown in Tables 1, 2 and 3, ReduMIS consistently achieves superior results compared to all other methods, with the exception of iSCO sometimes perform similarly, on ER, BA, and real-world graphs.
Additionally, although learning-based algorithms are claimed to be more efficient, they require significant training time and GPU memory, which is over our resource constraint for graphs with more than 3000 nodes, while ReduMIS can handle graphs with up to $1\times 10^{6}$ nodes (See Table 7). Although iSCO performs close to ReduMIS, it requires significant GPU memory and we are unable to get results for dense graphs with 10000 nodes. Its performance also become worse than ReduMIS for larger graphs.
@@
[…]
Heuristics GPU-acc Learning-based $n$ $m$ Deg-Greedy OnlineMIS ReduMIS iSCO PCQO LwD LTFT DIFUSCO DiffUCO 100 5 39.50 39.50 39.50 39.50 33.88 39.50 38.12 39.38 39.25 15 21.00 21.63 21.63 21.62 18.38 21.62 20.62 21.25 21.25 300 5 121.88 123.13 123.13 123.12 117.62 123.12 115.62 122.88 123.00 15 69.38 71.38 71.38 71.38 65.62 70.75 64.75 69.50 69.25 50 38.00 49.88 50.00 50.00 27.25 50.00 41.75 50.00 50.00 1000 5 412.63 417.13 417.13 417.12 410.88 416.00 385.75 417.12 416.00 15 232.75 245.00 246.38 246.25 231.00 243.12 224.25 236.38 240.75 50 107.50 115.75 116.88 116.75 107.88 113.00* 106.38 105.25 56.38† 150 81.50 150.00 150.00 150.00 109.62 150.00* 82.25 135.00† 3000 5 1236.63 1257.00 1257.13 1255.62 1238.75 1248.12 1177.25 1254.38 1252.63† 15 709.50 749.63 754.50 752.00 710.62 727.75* 661.88 726.38 737.38† 50 335.38 362.63 369.75 368.25 338.88 336.88* 323.88 142.25† 150 147.25 160.25 165.75 164.00 144.00 144.75 500 141.00 500.00 500.00 500.00 223.88 10000 5 4128.75 4206.00 4205.38 4190.47 4161.88 15 2377.88 2525.00 2534.00 2517.07 2428.13 50 1122.13 1228.88 1251.13 1241.98 1184.63 150 511.25 555.75 580.13 573.37 537.63 500 192.50 209.13 217.50 198.75 1500 67.25 1500.00 1500.00 1500.00
@@
While ReduMIS consistently outperforms AI-based methods in most cases, the performance gap is small on small or sparse graphs. On ER graphs when $n=100,300$ and average degree $d=10$, LwD has results within $1\%$ gap from ReduMIS. On BA graphs for $n=300,1000$ and $m=5$, LwD, DIFUSCO, and DiffUCO has less than $0.5\%$ gap from ReduMIS. Across real-world datasets all tested algorithms produce independent sets of similar sizes on average. However, as the graph becomes larger or denser, the performance gap between ReduMIS and AI-based algorithms enlarges. On ER graphs, when $n=1000$, there is a clear performance gap between AI-based algorithms and ReduMIS, with the only exception of the sampling based iSCO. When $d=10$, DiffUCO and LwD still reaches 98% of ReduMIS’s performance. For denser graphs($d=100$), DiffUCO and LwD only reaches 96% of ReduMIS’s performance. This gap widens further for $n=3000$, where AI-based algorithms perform significantly worse than ReduMIS and sometimes fail to outperform simple heuristic Deg-Greedy. A similar trend has also been observed on BA graphs. Classical solvers have no difficulty handling graphs with a million edges, but learning-based implementations struggle to scale up to that size.

### v2

@@ Abstract
This position paper argues that the machine learning community should fundamentally rethink how AI-inspired methods are developed and evaluated for combinatorial optimization (CO). We present comprehensive empirical benchmarks comparing various recent AI-inspired GPU-based methods with several classical CPU-based methods solvers on the Maximum Independent Set (MIS) problem. Strikingly, even on in-distribution random graphs, leading AI-inspired methods are consistently outperformed by state-of-art classical solver KaMIS, and some AI-inspired methods frequently fail to surpass even the simplest degree-based greedy heuristic. To better understand the source of these failures, we introduce a novel analysis, serialization, which reveals that non-backtracking AI methods, e.g. LTFT (which is based on GFlowNets), end up reasoning similarly to the simplest degree-based greedy, and thus worse than KaMIS.
Our findings reveal three core issues: (1) Limited benchmarks and evaluation—AI-inspired methods are often tested only on small instances with very limited inference time, which covers up issues with scalability and resource usage; (2) Intrinsic hardness and learning limits—even under ideal, in-distribution conditions, learning-based approaches lag behind classical heuristics, highlighting inherent barriers that receive little attention; and (3) Insufficient use and understanding of classical heuristics—current learning frameworks often neglect to incorporate effective classical techniques.
Although we use MIS as testbed, similar gaps and challenges have been reported in other combinatorial optimization problems, suggesting broader relevance for our recommendations. We propose that future research must address these issues by rigorous benchmarking, deepening theoretical and empirical understanding of learning limitations, and integrating classical heuristics into AI-inspired methods in a principled, evidence-driven manner.
@@ 1 Introduction
Combinatorial optimization (CO) lies at the core of numerous scientific and engineering studies, encompassing applications in network design, resource allocation, healthcare, and supply chain (du1998handbook, ; hoffman2000combinatorial, ; zhong2021preface, ). Most CO problems require selecting an optimal solution from a discrete but often exponentially large set of candidates and are typically NP-hard (papadimitriou1998combinatorial, ). While exact algorithms and advanced heuristics have long provided high-quality solutions in practice (e.g., Gurobi (gurobi, )), the last decade has seen a surge in machine learning (ML) and AI-inspired approaches to CO. These approaches—most notably, graph neural networks (GNNs), reinforcement learning (RL), and hybrid paradigms—promise to learn effective heuristics or even full solvers directly from data (li2018combinatorial, ; ahn2020learning, ; kool2018attention, ; zhang2021solving, ; bengio2021machine, ; cappart2023combinatorial, ).
Early successes—such as pointer networks and RL-based methods for routing problems (kool2018attention, ), GNN-based frameworks for MIS and clique (khalil2017learning, ; li2018combinatorial, ), and hybrid algorithms combining classical solvers with learned advice (balcan2018learning, ; nair2020solving, )—fueled excitement that ML-driven solvers might soon rival or surpass specialized algorithms across a broad range of CO tasks. More recent AI-inspired algorithms, such as diffusion models (sun2023difusco, ; sanokowskidiffusion, ), GPU-accelerated sampling (sun2023revisiting, ), direct optimization (alkhouri2024dataless, ), and -general-purpose tools like GFlowNets (bengio2021flow, ; bengio2023gflownet, ; zhang2023let, ), have been claimed to continue these improvements.
However, over the years, doubts have been raised about the practical gains of AI-inspired approaches. Empirical and theoretical studies raised concerns about their ability to generalize, to scale, and truly outperform strong classical baselines. A few years ago, (angelini2023modern, ) found that a GNN-based MIS algorithm (schuetz2022combinatorial, ) failed to outperform simple greedy baselines; (boether_dltreesearch_2022, ) showed that replacing heuristic components in tree search for MIS with GNN guidance offered little to no benefit, as performance largely stemmed from embedded classical reductions. Other works point to fundamental theoretical barriers for GNNs in CO (gamarnik2023barriers, ), and ML-based approaches for problems such as TSP have also struggled (joshi2021learning, ; xia2024position, ). These findings collectively suggested that it is difficult to replace classical heuristics with end-to-end learning and AI methods. Our paper evaluates more recent methods and our results suggest the same conclusion continues to hold.
This paper explores similar issues using the latest AI methods. We choose in our comparative study on a single NP-complete problem, Maximum Independent Set (MIS) (“given an undirected graph find the largest subset of nodes with no edges present between any pair within the subset”). The reason for this choice are as follows. The MIS problem’s simple formulation has tempted researchers to try a range of solution strategies, from simple greedy heuristics to sophisticated heuristic solvers like KaMIS (lamm2017finding, ; hespe2019scalable, ). The MIS problem is interesting and nontrivial even on random graphs, and this allows a fair testing ground for AI-based methods, whose train/test paradigm is most directly suited to settings with a clean distribution on inputs. Unsurprisingly, AI-inspired methods of all sorts have been applied to MIS, including GNN (li2018combinatorial, ), RL (ahn2020learning, ; zhang2023let, ), and diffusion models (sun2023difusco, ; sanokowskidiffusion, ). We evaluate (Section 3) a broad suite of AI-inspired and classical MIS solvers on a variety of graph types, with an emphasis on in-distribution random graphs. The results (Section 4) are revealing: even on test graphs drawn from the same distribution as the ML models’ training data, the state-of-the-art classical algorithm KaMIS, which uses far less computational resources while finding larger independent sets. Even more surprisingly, some AI-inspired methods frequently fail to outperform even the simplest degree-based greedy heuristic. These findings demonstrate that current AI-inspired methods for MIS have yet to replicate, let alone surpass, the effectiveness of well-established heuristics -—sometimes not even the simplest one.
The ML and AI community should rethink its approach to combinatorial optimization, focusing on the following issues (details in Section 6):
(1) Limited benchmarks and unrealistic evaluation: Current research often evaluates AI-inspired methods on small, quick-inference instances, failing to capture scalability and resource usage. Future work should establish rigorous, realistic benchmarks and report computational costs.
(2) Intrinsic hardness and learning limits: Even with well-defined distribution and entirely in-distribution training data, learning-based methods lag behind classical heuristics, indicating a need for deeper theoretical and empirical analysis of the limitations of learning algorithms.
(3) Insufficient use and understanding of classical heuristics: Progress will require principled integration of effective classical heuristics into AI-inspired frameworks, informed by careful analysis.
Although our testbed is MIS, we believe these positions broadly apply to other CO problems. Similar performance gaps and benchmarking issues have been found in TSP (joshi2021learning, ; sun2023difusco, ; xia2024position, ), and recent theoretical work highlights barriers for GNN-based methods across CO tasks (gamarnik2023barriers, ).
To further support our position that AI-based methods are not providing novel insight into MIS problem, Section 5 presents ablation studies to understand why some AI-inspired methods fail to surpass the simplest Deg-Greedy method. We introduce a new mode of analysis, serialization, which reveals structural similarities between certain AI-driven methods and basic greedy heuristics (Sections 5.1 and 5.2). We also explore whether AI-inspired methods can improve their solutions through local search post-processing (Section 5.3), but find limited improvements. Our findings underscore both the intrinsic difficulty of MIS for learning-based approaches and the enduring value of classical heuristics. We conclude by outlining issues and future directions in Section 6, and also provide alternative views in Section 7.
2 Related Works
Classical and heuristic methods for MIS:
Classical methods for MIS range from simple greedy algorithms to advanced solvers like KaMIS which involves a number of heuristics. There are various existing heuristics for MIS, such as reduction techniques (butenko2002finding, ; xiao2013confining, ; akiba2016branch, ), local search (andrade2012fast, ), and evolutionary algorithms (back1994evolutionary, ; borisovsky2003experimental, ; lamm2015graph, ). KaMIS (lamm2017finding, ; dahlum2016accelerating, ; hespe2019scalable, ) was developed based on these heuristics. In addition, MIS can be formulated into a binary integer programming problem (nemhauser1975vertex, ), which can be solved by the state-of-art integer programming solver Gurobi gurobi .
Machine learning for combinatorial optimization:
In recent years, various ML-based algorithms have been developed for the MIS problem. Most of them are based on graph neural networks (GNNs). Some use supervised learning (li2018combinatorial, ; sun2023difusco, ; li2024distribution, ) and requires labelling training data using classical solvers. Alternatively, those based on reinforcement learning (khalil2017learning, ; ahn2020learning, ; qiu2022dimes, ; sanokowski2023variational, ) and other unsupervised learning objectives (karalias2020erdos, ; sun2022annealed, ; zhang2023let, ; sanokowskidiffusion, ) do not require labeled training data. (zhang2023let, ) (LTFT) uses GFlowNets (bengio2021flow, ) which is related to reinforcement learning. Notably, (ahn2020learning, ; sanokowski2023variational, ; zhang2023let, ) model MIS problem as a Markov Decision Process (MDP) and generate the solution step-by-step (autoregressively). While (sanokowski2023variational, ) fixes the order of node updates, (ahn2020learning, ) (LwD) and (zhang2023let, ) (LTFT) choose which node to update at each step so that they have a “natural" or “pseudo-natural" serialization and most suitable for our analysis in Section 5.1 and Section C.5. (sun2023difusco, ) (DIFUSCO) and (sanokowskidiffusion, ) (DiffUCO) both use diffusion model.
Most of the algorithms above also work on other types of graph CO problems, such as Maximum Cut, Maximum Vetex Cover, and Minimum Dominating Set. Some (khalil2017learning, ; qiu2022dimes, ; sun2023difusco, ) also work for the Travelling Salesman Problem (TSP).
Other GPU-based solvers for CO:
Recently, some non-learning GPU-based solvers for CO problems have been developed. (sun2023revisiting, ) developed a GPU-accelerated sampling based method which works on MIS, Max Cut, and TSP. (schuetz2022combinatorial, ; ichikawa2023controlling, ) uses GNNs conduct non-convex optimization for MIS without machine learning. (alkhouri2024dataless, ) uses direct quadratic optimization without GNNs for MIS.
Benchmarks for MIS
(boether_dltreesearch_2022, ) provides a benchmark for several MIS algorithms including Gurobi (gurobi, ), KaMIS (ReduMIS(lamm2017finding, ), Intel-Treesearch (li2018combinatorial, ), DGL-Treesearch, and LwD ahn2020learning . It is the only MIS benchmark we know about including recent AI-inspired method, though it only focuses its comparsion for learning-based tree search algorithms. This benchmark covers various types of random graphs and several real-world datasets, so it is a good reference benchmark for comparison over different types of graphs. Though unlike our benchmark, it does not provide comparison across various size and density for random graphs. Our benchmark fills this gap and provides a more detailed comparison. We also include many newer AI-inspired algorithms, and greedy algorithms which leads to detailed analysis like serialization. (angelini2023modern, ) provides a comparison between (schuetz2022combinatorial, ) and Deg-Greedy over regular graphs. The algorithm papers also report experiments comparing with previous algorithms, but those comparisons usually only focus on a few datasets and a few selected baselines.
3 Benchmarking AI-inspired Methods for Maximum Independent Set
@@ 3.2 MIS algorithms
Deg-Greedy (Degree-based greedy) Simplest heuristic: always picks a node with the smallest degree in the current graph, add to the current independent set, and delete that node and all of its neighbors from the graph. Most papers on AI-inspired methods do not compare with this baseline.
OnlineMIS and ReduMIS are two variants of the MIS solver KaMIS, mainly consists of three alternating steps: greedy, local search, and graph reductions. OnlineMIS (dahlum2016accelerating, ) only applies a simple reduction after local search, while ReduMIS lamm2017finding applies many graph reduction techniques.
iSCO (sun2023revisiting, ) is a GPU-accelerated sampling-based method, incorporating gradient-based discrete MCMC and simulated annealing. The MCMC is designed based on the Metropolis-Hasting algorithm, which if given enough time (exponential), can get the optimal solution.
PCQO (alkhouri2024dataless, ) directly optimizes the quadratic loss function of the MIS using gradient descent. It is sensitive to optimization hyperparameters, so hyperparameter search is required for achieving good results. Extensive hyperparameter search may give better results than in our experiments.
LwD (ahn2020learning, ) is a reinforcement learning based algorithm which models the problem as a Markov Decision Process (MDP) and requires a dataset (without solutions) to train the policy. In each step, several nodes are added to the independent set and are never deleted. We call it a non-backtracking algorithm.
LTFT (zhang2023let, ) is also a non-backtracking MDP-based algorithm similar to LwD, but it only selects one node at a time, which is decided by GFlowNets (bengio2021flow, ). Thus, it has a very similar procedure to Deg-Greedy. The algorithm requires a dataset (without solutions) to train the neural network.
DIFUSCO and DiffUCO DIFUSCO (sun2023difusco, ) is an end-to-end one-shot MIS solver using diffusion models and requires a supervised training dataset with solutions. DiffUCO (sanokowskidiffusion, ) also uses diffusion model but with unsupervised learning and annealing techniques. It requires a training dataset without solutions.
@@
3.3 Datasets and experiment details
We generated and experimented on 2 random graph types: Erdős-Renyí (ER) graphs (erdos59a, ) and Barabási–Albert (BA) graphs (albert2002statistical, ), across various sizes and densitites. We also experimented on existing datasets: synthetic RB graphs (xu2000exact, ) from (zhang2023let, ), and real-world datasets form (yanardag2015deep, ). Due to space limit, we only present ER graphs in main text. Descriptions of other graphs are in Section B.2.
Erdős-Renyí (ER) graph
(erdos59a, ) are random graphs where edges are connected uniformly at random (with a fixed probability or number of edges). We vary $2$ parameters for ER graphs, number of nodes $n$ and average degree $d$, by fixing number of edges at $\frac{nd}{2}$. Previous works (boether_dltreesearch_2022, ; ahn2020learning, ; sun2023difusco, ; zhang2023let, ; alkhouri2024dataless, ) used it as test graphs for MIS, though without varying parameters as we did.
We set $24$-hr time limit for KaMIS (OnlineMIS and ReduMIS) since it runs on a single CPU thread with 32GB memory, and our benchmark focuses on performance on solution quality. ReduMIS runs in less than one hour on small graphs $(n\leq 3000)$. Given 24 hours ReduMIS can handle much larger graphs (up to $n\approx 1e6$). For iSCO and learning-based algorithms, we report results within our computational limit (generally a single 80GB A100 for $96$hrs, details in Appendix B.) We test PCQO for $n\leq 10000$ because the performance degrades quickly for large graphs using default hyperparameter search domain.
4 Performance Gap: Classical Methods Outperform AI-inspired Methods
In this section, we present our main experiment results. The results for various algorithms on Erdős-Renyí (ER) graphs are in Table 1. Due to space limit, the results for other graphs are in appendix: Barabási–Albert (BA) graphs in Table 5, RB graphs (xu2000exact, ) in Table 6, and real-world graphs in Table 7.
@@ 4 Performance Gap: Classical Methods Outperform AI-inspired Methods
CPU-based GPU-acc Learning-based $n$ $d$ Deg-Greedy OnlineMIS ReduMIS iSCO PCQO LwD LTFT DIFUSCO DiffUCO 100 10 29.25 30.50 30.50 30.62 30.63 30.38 28.62 30.25 30.02 30 13.63 14.00 14.75 14.50 14.50 14.38 13.12 13.88 13.92 300 10 77.50 93.88 94.38 94.75 94.63 94.25 88.62 93.50 91.84 30 44.50 47.88 47.88 47.62 47.63 46.88 43.25 43.88 45.04 100 16.13 18.00 18.38 18.00 18.00 17.00 16.25 16.62 16.93 1000 10 303.25 314.75 316.13 315.62 310.00 311.25 297.00 303.88 311.67 30 151.00 158.88 163.75 163.50 158.63 158.38 150.00 143.75 152.55† 100 60.63 64.75 66.63 66.50 60.13 63.88 60.88 55.38 63.55† 300 22.25 25.00 25.75 24.62 23.25 19.12* 22.62 20.88 18.36† 3000 10 907.13 947.25 954.25 950.88 923.25 934.12 888.25 902.00 935.28† 30 451.88 480.88 493.13 491.62 464.25 473.25 449.00 413.38 459.57† 100 183.63 194.38 201.50 200.38 185.63 190.75* 184.00 171.38 194.07† 300 73.50 77.63 80.75 78.88 69.25 73.88 1000 23.38 26.00 26.25 23.00 23.62 10000 10 2999.88 3161.88 3173.62 3149.92 2569.75 30 1498.00 1607.50 1639.88 1625.47 1331.63 100 613.75 650.00 670.88 543.13 300 249.00 258.38 272.25 218.63 1000 87.13 91.50 94.25 77.50 3000 29.63 33.00 33.25 26.63
@@
Our first main finding is that, current AI-inspired algorithms do not outperform the best classical heuristic ReduMIS in terms of performance. As shown in Tables 1, 5 and 7, ReduMIS consistently achieves superior results compared to all other methods, with the exception of iSCO sometimes perform similarly.
Although learning-based algorithms are claimed to be more efficient, they require significant training time and GPU memory, which is over our resource constraint for graphs with more than 3000 nodes, while ReduMIS can handle graphs with up to $1\times 10^{6}$ nodes (See Table 8). Although iSCO performs close to ReduMIS, it requires significant GPU memory and we are unable to get results for dense graphs with 10000 nodes. Its performance also become worse than ReduMIS for larger graphs.
@@
While ReduMIS consistently outperforms AI-based methods in most cases, the performance gap is small on small or sparse graphs. On ER graphs when $n=100,300$ and average degree $d=10$, LwD has results within $1\%$ gap from ReduMIS. However, as the graph becomes larger or denser, the performance gap between ReduMIS and AI-based algorithms enlarges. On ER graphs, when $n=1000$, there is a clear performance gap between AI-based algorithms and ReduMIS, with the only exception of the sampling based iSCO. When $d=10$, DiffUCO and LwD still reaches 98% of ReduMIS’s performance. For denser graphs($d=100$), DiffUCO and LwD only reaches 96% of ReduMIS’s performance. This gap widens further for $n=3000$, where AI-based algorithms perform significantly worse than ReduMIS and sometimes fail to outperform simple heuristic Deg-Greedy. Classical solvers have no difficulty handling graphs with a million edges, but learning-based implementations struggle to scale up to that size.